OpenStudy (wpeyron99):

My shoes were only $19.50 but I also had to pay the 14% tax. What was my total? help plz

OpenStudy (helpppp):

turn your percent to a decimal.

OpenStudy (helpppp):

Multiply it by the original price. And add the result to that original

OpenStudy (wpeyron99):

so the the percent will look like 0.14

OpenStudy (sshayer):

\[19.50+(19.50 \times \frac{ 14 }{ 100 })=?\]

OpenStudy (helpppp):

Yes, 0.14 would be the decimal. multiply it by your original.

OpenStudy (wpeyron99):

so is the answer $ 27.30

OpenStudy (sshayer):

check your calculations again.

OpenStudy (wpeyron99):

kk

OpenStudy (wpeyron99):

is it $78

OpenStudy (kittiwitti1):

let's leave off the signs just to minimize confusion. original price: 19.50 tax: 14% to get total price you must add tax, which is a percentage of your original price, to the original price. this means: 14% (19.50) + 19.50

OpenStudy (kittiwitti1):

14% of 19.50 is also 0.14\(\times\)19.50

OpenStudy (kittiwitti1):

add this \(\Large{↑}\) to the original price of 19.50

OpenStudy (helpppp):

Dats what I was sayin :P

OpenStudy (sshayer):

19.50*14%=2.73 19.50+2.73=?

OpenStudy (wpeyron99):

thanx

OpenStudy (sshayer):

yw